OpenStudy (anonymous):

write a rule for the -nth term of the arithmetic sequence. then find the 12th term. 1)24,35,46 .... 2) 31,17,3 ......

OpenStudy (anonymous):

@jim_thompson5910 can you explain this to me please

jimthompson5910 (jim_thompson5910):

24,35,46 what's the first term of the sequence above

OpenStudy (anonymous):

like the first number?...24 ?

jimthompson5910 (jim_thompson5910):

yep good

jimthompson5910 (jim_thompson5910):

what is the common ratio of 24,35,46, ...

OpenStudy (anonymous):

keep adding 11

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

first term: a1 = 24 common ratio: d = 11

jimthompson5910 (jim_thompson5910):

now plug these two things into the formula below a_n = a1 + d(n-1) and simplify

OpenStudy (anonymous):

what do they mean by 12th term? how?

jimthompson5910 (jim_thompson5910):

what did you get for the nth term

OpenStudy (anonymous):

for the n1 ...do I just leave it like that?

OpenStudy (anonymous):

im at an=24+11(n1)

jimthompson5910 (jim_thompson5910):

a_n = a1 + d(n-1) a_n = 24 + 11(n-1) a_n = 24 + 11n-11 ... distribute now combine like terms

OpenStudy (anonymous):

do i add 11?

OpenStudy (anonymous):

or just make it the n stand alone

jimthompson5910 (jim_thompson5910):

24 - 11 = ??

OpenStudy (anonymous):

13

jimthompson5910 (jim_thompson5910):

so the nth term is a_n = 11n + 13

jimthompson5910 (jim_thompson5910):

to find the 12th term, plug in n = 12 and evaluate/simplify

OpenStudy (anonymous):

to where?

jimthompson5910 (jim_thompson5910):

a_n = 11n + 13 a12 = 11*12 + 13 ...plug in n = 12 a12 = ??

OpenStudy (anonymous):

145

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's the 12th term of 24,35,46, ...

OpenStudy (anonymous):

the next one ... would i use the same formula ? an=a1+d(n-1)

jimthompson5910 (jim_thompson5910):

correct since that's also an arithmetic sequence

OpenStudy (anonymous):

and for example on the next question it says find the tenth term of the arithmentic sequence in which a1=-8 and d=-2 ....how would i find that?

jimthompson5910 (jim_thompson5910):

first find the nth term, then plug in n = 10 to find the tenth term

OpenStudy (anonymous):

ok got it ! how would i answer the question . a theater has 13 seats in the first row. each row after the first has 4 more seats than the previous row. write a rule for the -nth term of this situation ?

jimthompson5910 (jim_thompson5910):

hint: first term: 13 common difference: 4

OpenStudy (anonymous):

how many seats are in the 11th row?

OpenStudy (anonymous):

how would i find that

jimthompson5910 (jim_thompson5910):

plug in n = 11 and evaluate

OpenStudy (anonymous):

plug it in to what number

jimthompson5910 (jim_thompson5910):

to what you got for your nth term

jimthompson5910 (jim_thompson5910):

what's your nth term

OpenStudy (anonymous):

i got an=4n+12

OpenStudy (anonymous):

would it be 56?

jimthompson5910 (jim_thompson5910):

nth term an = 4n + 9 plug in n = 11 to get the 11th term a11 = 4*11+9 = 53